# Efficiently Serving Large Multimodal Models Using EPD Disaggregation

### v2

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead. This step negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput (E2ETP). We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources. Unlike current systems, which bundle encoding and prefill together, our approach decouple these steps unlocking new opportunities and optimizations. These include a new mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize encoding load within a request, a module to find the optimal resource allocation for disaggregated serving, and a novel role switching method to handle changing workload characteristics. Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ less utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images/request, and 2.2$\times$ larger KV caches. Further, it leads to significant improvements in latency metrics (TTFT up to 71% reduction) and end-to-end throughput (up to 57% reduction), compared to systems that do not disaggregate.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized language understanding and reasoning, achieving superhuman performance across a variety of tasks [1, 4]. Recently, the scope of these models has expanded to include multiple modalities, such as images, audio, and videos, resulting in the advent of Large Multimodal Models (LMMs) [19, 11, 5]. LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries.
Figure 1: Above: Aggregated system, where E and LLM are coupled and must run on the same GPU. Below: Disaggregated system, where E runs on the 1st GPUs and LLM runs on the rest of three GPUs. Multimodal encoder and LLM are denoted by E and LLM, respectively.
@@
However, serving LMMs in an efficient manner presents unique challenges. Meeting strict Service Level Objectives (SLOs) such as end-to-end throughput (E2ETP), time to first token (TTFT), and time per output token (TPOT) becomes increasingly difficult given the additional computational and memory demands introduced by processing multimodal data [12]. Unlike LLMs, where inference involves prefill and decoding stages, LMMs require an additional encoding stage to process raw multimodal inputs (e.g., images or videos) into tokenized representations. This stage is computationally intensive, especially for high-resolution or complex multimodal inputs, and often produces a substantial number of additional tokens [17]. The resulting increase in tokens inflates resource consumption and leads to quadratic growth in prefill-stage compute demands, adversely impacting SLO attainment.
@@
Disaggregating the prefill stage from the decode stage has emerged as a well-studied solution in the literature to improve inference efficiency for LLMs [21, 15, 14, 7, 6]. By assigning separate resources to each stage, prefill-decode disaggregation enables independent optimization of batching, scheduling, and resource allocation strategies, significantly enhancing system throughput and memory utilization. However, these techniques fail to address the challenges of LMM deployment, where the addition of an encoding stage fundamentally alters the resource dynamics. The encoding stage, with its high computational and memory overhead, introduces token inflation and dependency bottlenecks that ripple across subsequent stages, necessitating a rethinking of disaggregation strategies to accommodate multimodal workloads.
Figure 2: Impact of disaggregation on batch size and number of images/ request for the MiniCPM-V 2.6 model
@@
These challenges of LLM deployment present opportunities for optimization that current serving systems do not consider. Currently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs. Sequential execution of encoding and prefilling introduces interference. This interference is demonstrated graphically in Figure 1. In the aggregated approach (Figure 1 top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$). As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, revealing that current solutions are inadequate for LMM workloads. By disaggregating encode from prefill, the system may be optimized to reduce interferences for specific workloads. From another perspective, simultaneous loading of both the Multimodal encoder (MME) and LLM onto the same GPUs in aggregated setups restricts memory for processing multiple high resolution images or larger batch sizes. In preliminary investigations, by removing the LLM and associated KV cache, we observe a significant increase in maximum batch sizes and supported number of images/ request (Figure 2).
@@
Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each with dedicated resources to operate independently. This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention. As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
@@
  • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages. Our approach addresses key challenges of inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
  • We propose intra-request parallelization (IPR) that shards a request into independent encoding jobs which can be executed in parallel, significantly reducing first token latency.
  • We formulate the resource allocation problem as an optimization over batch sizes, scheduling strategies, and parallelization approaches for each pipeline stage. Further, we provide a simple black box optimizer that uses workload samples to approximate the optimal config for a given workload.
  • We develop dynamic role switching capability that monitors the system for any bottlenecks and allows efficient reallocation of resources between E, P and D stages by efficiently switching ($\sim$0.7s) the role of an instance among these stages. This ensures system can respond to any changes in workload requirements during online serving.
  • We conduct evaluations on various popular LMM models, namely MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B using real-world and synthetic workloads. The results demonstrate the superiority of our method across various aspects. For instance, our approach requires up to 15x less memory, facilitating up to 22x larger batch sizes, 10x more images/request, and 2.2x larger KV caches. Further, our method significantly improves performance in terms of SLO attainment (up to 90–100% more) and TTFT latency (up to 71% reduction).
@@ 2 Related Work
In this section, we describe feasible approaches in the literature for multi-modality model serving and their limitations and then review the use of the disaggregation technique.
Multi-modality Model Serving. There are two ways to enable a multi-modality model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM [9], SARATHI [2], and Orca [20]; and (2) leveraging the open-sourced code from recent works of improving multi-modality model inference. The first way fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multi-media data. In the second way, recent advancements such as KV cache eviction [10, 13] and compression [8] focus on specific features and limited scenarios. For instance, Inf-MLLM [13] enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios. While their implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios. To the best of our knowledge, we are the first to propose an LMM serving system that leverages disaggregation and a series of integrated techniques to enable better resource allocation and improved SLOs.
Disaggregated Serving. Disaggregated serving has emerged as a promising technique in large model serving. By decoupling the prefill and decode stages, systems like SplitWise [14], DistServe [21], and DéjàVu [16] mitigate interference between these phases, enabling finer control over Time to First Token (TTFT) and Time Per Output Token (TPOT). However, these systems primarily target LLMs and overlook the encoding step required for LMMs, which is tightly coupled with prefill. Recent work, such as Mooncake [15] and PD-Serve [7], has extended disaggregation to include KV cache management and other system-level optimizations. Despite these advancements, they remain limited in their applicability to LMMs. Disaggregating the encoding phase from prefill could unlock new opportunities for optimizing LMM serving. For instance, when handling requests with long videos, encoding the frames could be executed in parallel, reducing head-of-line blocking and improving overall system efficiency. This approach would enable more flexible batching and scheduling strategies, ultimately achieving better latency and throughput for multimodal workloads.
@@ 3 Method
In this section, first, we describe the EPD disaggregation framework (Sec. 3.1), followed by a detailed explanation of the system design and optimizations (Sec. 3.2), that includes the asynchronous token transfer, intra-request parallel, optimized resource allocation, and dynamic role switching.
@@ 3.1 EPD Disaggregation
As shown in Figure 3, disaggregating an LMM system involves breaking the inference process into multiple stages: encoding, prefill, and decode. Transitions between these stages—EP-migration and PD-migration—handle the transfer of data from encoding to prefill and from prefill to decode, respectively. Formally, we denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the text output as $o$, and provide details for each step and transition as follows:
Encoding (E): The multimodal input $i_{m}$ is processed by the MME $E$, converting it into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs. These tokens form a high-dimensional multimodal embedding for the next pipeline stage. This process can be denoted as $v_{t}^{e}=E(i_{m})$.
EP-migration: Once the $E$ stage completes processing, the generated multimodal tokens $v_{t}^{e}$ are transferred to the prefill stage $P$. This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache. This process is expressed as $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$. Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
Prefill (P): The multimodal tokens $v_{t}^{p}$ arrive at the Prefill stage $P$, along with the text prompt $i_{p}$. In this stage, the system initializes the Key-Value (KV) cache $kv_{1}^{p}$ and generates the first output token $o_{1}^{p}$. The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage. The process is denoted by $kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$.
PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$. This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker. This is denoted as $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
Decode (D): The decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the token generated in the previous iteration $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$. This autoregressive process continues until the desired output sequence is generated, as denoted by $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
@@ 3.2 System Design and Optimization
Cache transfers occur asynchronously when an instance pulls a request from its queue, ensuring the transfer occurs once the downstream instance is ready. Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on ensuring smooth token transfers between stages (Section 3.2.1), reducing latency (Section 3.2.2), managing resources effectively (Section 3.2.3), and adapting to changing workloads (Section 3.2.4). The ablation of these features are shown in Sec. 4.4.
@@ 3.2.2 Intra-Request Parallel (IRP)
Multimodal requests often include multiple high-resolution images in many practical scenarions like autonomous driving and video question answering. In modern LMMs, these images are further converted into a large number of patches that needs to be processed through a computationally-heavy MME. This significantly increases the computation load, making the encoding process a major bottleneck.
@@ 3.2.3 Optimized Resource Allocation
Let $f(\cdot)$ represent the system’s performance metric (e.g., goodput, see Section 4). Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe [21]—to evaluate performance metrics efficiently. The objective is to maximize performance while minimizing GPU usage, which inherently reduces pipeline inefficiencies (e.g., idle time) and improves resource utilization. Formally, we solve
@@
Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix A for details). We use Bayesian optimization [3] to solve Problem 1.
@@ 4 Experiments
In this section, we analyze and compare the performance of the proposed EPD disaggregation method against various baselines. We start with an end-to-end generation performance analysis in Sec. 4.1, followed by an examination of first token latency in Sec. 4.2. We then evaluate the memory savings achieved by EPD in Sec. 4.3. Finally, an ablation analysis of key system components is presented in Sec. 4.4. Additional experiments are detailed in the Appendix, including 1) a throughput comparison in offline settings (Appendix C); 2) an extension of our framework to Neural Processing Units (NPUs) demonstrating a 10% performance improvement (LABEL:sec:appendix-npu), and 3) additional SLO and memory experiments (Appendix D).
@@
Figure 5: SLO attainment for end-to-end inference across various models and image counts per request. Columns (a), (b), and (c) correspond to the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively. The top and bottom rows are with 2 and 4 images per request, respectively. As observed, EPD significantly outperforms all baselines across all configurations.
Baselines: We compared the proposed EPD method against two popular baselines: DistServe [21] and vLLM [9]. The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs. Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying the block manager to accommodate multimodal tokens. The vLLM baseline adopts a monolithic architecture, where all three stages run on the same set of GPUs.
Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [19], InternVL2-8B, and InternVL2-26B [5]. These LMMs are renowned for their advanced capabilities in processing and understanding multimodal data. Detailed descriptions of the LMMs and their sizes can be found in Appendix B.2.
Datasets: To evaluate performance across diverse scenarios, we conducted experiments using two datasets. The first was a synthetic workload, that allowed setting custom parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings. Unless otherwise specified in our experiments, the input prompt length was set to 22 tokens. The second dataset was NextQA [18], a benchmark video question-answering dataset. Unlike synthetic workload, NextQA features human-annotated questions and answers, providing a more realistic reflection of real-world video request distributions.
Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The related performance metrics are listed below.
@@
  • TPOT: The average time interval between each output token (excluding the first token).
@@
  • Goodput: The hightest request rate at which 90% or more SLO attainment is achieved.
For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory. Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes. (See section 4.3)

### v3

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead. This step negatively affects key Service Level Objectives (SLOs), such as time to first token (TTFT) and time per output token (TPOT). We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources. Unlike current systems, which bundle encoding and prefill together, our approach decouples these steps, unlocking new opportunities and optimizations. These include a mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize the encoding load within a request, a module for optimal resource allocation for disaggregated serving, and a novel role-switching method to handle changing workload characteristics. Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ lower peak memory utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images per request, and 2.2$\times$ larger KV caches. Furthermore, it leads to significant improvements in SLO attainment (up to 90–100% improvement) and TTFT (up to 71% reduction), compared to systems that do not disaggregate. The code is available at https://github.com/vbdi/epdserve.
Efficient inference, Disaggregated inference, Multimodal modal inference, LLM inference
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized language understanding and reasoning, achieving superhuman performance on a variety of tasks (Achiam et al., 2023; Chang et al., 2024). Recently, the scope of these models has expanded to include multiple modalities, such as images, audio, and videos, leading to the emergence of Large Multimodal Models (LMMs) (Yao et al., 2024; Liu et al., 2023; Chen et al., 2024). LMMs enable users to interact with diverse data types, such as posing questions about visual scenes or analyzing audio clips, thereby unlocking novel applications across fields like healthcare, autonomous systems, and creative industries (Luo et al., 2025).
Figure 1: Aggregated (top) vs. disaggregated (bottom) system architectures. In the aggregated setup, the encoder (E) and LLM share the same GPUs, leading to interference between encode and prefill stages (e.g., LLM-4 delays E5). Disaggregation isolates these stages across GPUs, reducing contention and improving utilization. This separation enables better performance under multimodal workloads, addressing key limitations of existing systems.
@@
However, serving LMMs in an efficient manner presents unique challenges. Meeting strict Service Level Objectives (SLOs), such as time to first token (TTFT) and time per output token (TPOT), becomes increasingly difficult given the added computational and memory demands of processing multimodal data (Alvar et al., 2025; Liu et al., 2024). Unlike LLMs, where inference involves prefill and decoding stages, LMMs require an additional encoding stage to process raw multimodal inputs (e.g., images or videos) into tokenized representations. This stage is computationally intensive, especially for high-resolution or complex multimodal inputs, and often produces a substantial number of additional tokens (Wu et al., 2023). The resulting token inflation increases resource consumption and leads to quadratic growth in prefill-stage compute demands, adversely impacting SLO attainment.
@@
Disaggregating the prefill stage from the decode stage has emerged as a well-studied solution for improving LLM inference efficiency (Zhong et al., 2024; Qin et al., 2024; Patel et al., 2024; Jin et al., 2024; Hu et al., 2024). By assigning separate resources to each stage, prefill-decode disaggregation enables independent optimization of batching, scheduling, and resource allocation strategies, significantly enhancing system throughput and memory utilization. However, these techniques fall short in addressing LMM-specific challenges, as the addition of an encoding stage fundamentally changes the resource dynamics. The encoding stage adds significant compute and memory overhead, inflates tokens, and creates dependencies that affect later stages, requiring a fresh look at disaggregation strategies for multimodal workloads.
Figure 2: Impact of disaggregation on supported batch size and number of images per request for the MiniCPM-V 2.6 model. Removing the LLM from the GPU significantly increases capacity, enabling larger batches and higher-resolution inputs. This demonstrates the memory efficiency benefits of disaggregation.
@@
These challenges present optimization opportunities that current serving systems do not exploit. Presently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs. Sequential execution of encoding and prefilling introduces interference, as demonstrated in Figure 1. In the aggregated setup (top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$). As a result, such aggregation leads to suboptimal resource utilization and degraded SLO performance, revealing the inadequacy of current solutions for LMM workloads. Disaggregating encoding from prefill enables the system to reduce such interference under specific workloads. From another perspective, simultaneously loading both the multimodal encoder (MME) and the LLM onto the same GPUs in aggregated setups restricts the available memory for processing multiple high-resolution images or supporting larger batch sizes. In preliminary investigations, removing the LLM and its associated KV cache significantly increases the maximum batch size and the number of images per request (Figure 2).
@@
Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each stage dedicated resources to operate independently. This design enables customized strategies for batching, parallelization, and scheduling at each stage, optimizing resource utilization while reducing contention. As a result, EPD achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT and TPOT.
@@
  • We propose an efficient system for LMM inference that introduces the novel idea of disaggregating the encoding and prefill stages. Our approach addresses key challenges in inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
  • We introduce intra-request parallelization (IRP), which shards a request into independent encoding jobs that can be executed in parallel, significantly reducing first-token latency.
  • We formulate the resource allocation problem as an optimization over batch sizes, scheduling strategies, and parallelization approaches for each pipeline stage. Furthermore, we provide a simple black-box optimizer that uses workload samples to approximate the optimal configuration for a given workload.
  • We develop a dynamic role-switching capability that monitors the system for bottlenecks and enables flexible reallocation of resources across the E, P, and D stages by switching an instance’s role between stages. This ensures the system can respond effectively to changes in workload requirements during online serving.
  • We conduct evaluations on several popular LMM models—MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B—using both real-world and synthetic workloads. Results demonstrate the superiority of our approach: up to 15$\times$ lower memory usage, 22$\times$ larger batch sizes, 10$\times$ more images per request, up to a 90–100% improvement in SLO attainment, and up to 71% lower TTFT.
@@ 2 Related Work
In this section, we describe existing approaches in the literature for multimodal model serving and their limitations, and then we review the use of disaggregation techniques.
Multimodal Model Serving. There are two ways to enable a multimodal model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM (Kwon et al., 2023), SARATHI (Agrawal et al., 2023), and Orca (Yu et al., 2022); and (2) leveraging open-source code from recent works on improving multimodal model inference. The first approach fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multimedia data. In the second approach, recent advancements such as KV cache eviction (Li et al., 2024; Ning et al., 2024) and compression (Kang et al., 2024) focus on specific features and limited scenarios. For instance, Inf-MLLM (Ning et al., 2024) enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios. While these implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios. To the best of our knowledge, we are the first to propose an LMM serving system that leverages disaggregation and a series of integrated techniques to enable better resource allocation and improved SLOs.
Disaggregated Serving. Disaggregated serving has emerged as a promising technique in large model serving. By decoupling the prefill and decode stages, systems like SplitWise (Patel et al., 2024), DistServe (Zhong et al., 2024), and DéjàVu (Strati et al., 2024) mitigate interference between these phases, enabling finer control over TTFT and TPOT. However, these systems primarily target LLMs and overlook the encoding step required for LMMs, which is tightly coupled with prefill. Recent work, such as Mooncake (Qin et al., 2024) and PD-Serve (Jin et al., 2024), have extended disaggregation to include KV cache management and other system-level optimizations. Despite these advancements, they remain limited in their applicability to LMMs. Disaggregating the encoding phase from prefill could unlock new opportunities for optimizing LMM serving. For instance, when handling requests with long videos, the frames could be encoded in parallel, reducing head-of-line blocking and improving overall system efficiency. This approach would enable more flexible batching and scheduling strategies, achieving better latency and throughput for multimodal workloads.
@@ 3 Method
In this section, we first describe the EPD disaggregation framework (Section 3.1), followed by a detailed explanation of the system design and optimizations (Section 3.2), including asynchronous token transfer, intra-request parallelism, optimized resource allocation, and dynamic role switching.
@@ 3.1 EPD Disaggregation
As shown in Figure 3, disaggregating an LMM system involves dividing the inference process into three stages: encoding, prefill, and decode. Transitions between these stages—EP-migration and PD-migration—handle the transfer of data from encoding to prefill and from prefill to decode, respectively. We denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the output text as $o$. The steps are as follows:
Encoding (E): The multimodal input $i_{m}$ is processed by the multimodal encoder $E$, converting it into multimodal tokens $v_{t}^{e}=E(i_{m})$, which form a high-dimensional embedding for the next pipeline stage.
EP-migration: Once encoding completes, the generated tokens are transferred to the prefill stage $P$ via the EP-migration function $\psi_{EP}$, such that $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$.
Prefill (P): The prefill stage processes $v_{t}^{p}$ and the text prompt $i_{p}$ to produce the initial KV cache and first output token: $kv_{1}^{p},o_{1}^{p}=P(v_{t}^{p},i_{p})$.
PD-migration: The KV cache and token are passed to the decode stage via the PD-migration function $\psi_{PD}$, where $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
Decode (D): Decoding proceeds autoregressively, generating the next token $o_{t+1}^{d}$ and updating the cache: $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
This autoregressive process continues until the output sequence is fully generated.
@@ 3.2 System Design and Optimization
Cache transfers occur asynchronously when an instance pulls a request from its queue, ensuring the transfer occurs once the downstream instance is ready. Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on ensuring smooth token transfers between stages (Section 3.2.1), reducing latency (Section 3.2.2), managing resources effectively (Section 3.2.3), and adapting to changing workloads (Section 3.2.4). The ablation of these features is shown in Section 4.4.
@@ 3.2.2 Intra-Request Parallel (IRP)
Multimodal requests often include multiple high-resolution images in many practical scenarios like autonomous driving and video question answering. In modern LMMs, these images are further converted into a large number of patches that need to be processed through a computationally heavy MME. This significantly increases the computation load, making the encoding process a major bottleneck.
@@ 3.2.3 Optimized Resource Allocation
Let $f(\cdot)$ represent the system’s performance metric (e.g., goodput, see Section 4). Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe (Zhong et al., 2024)—to evaluate performance metrics efficiently. The objective is to maximize performance while minimizing GPU usage, which inherently reduces pipeline inefficiencies (e.g., idle time) and improves resource utilization. Formally, we solve
@@
Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix D for details). We use Bayesian optimization (Calvo et al., 2019) to solve Problem 1.
@@ 4 Experiments
In this section, we analyze and compare the performance of the proposed EPD disaggregation method against various baselines. We start with an end-to-end generation performance analysis in Section 4.1, followed by an examination of the first token latency in Section 4.2. We then evaluate the memory savings achieved by EPD in Section 4.3. Next, we present an ablation analysis of the key system components in Section 4.4. Finally, we present an extension of our framework to Neural Processing Units (NPUs) in Section 4.5. Additional experiments are detailed in the Appendix, including 1) a throughput comparison for offline and heterogeneous settings in Section A.3, 2) additional analysis of the implementation on NPUs in Appendix F, 3) an experiment extending our framework to the audio modality in Section A.1, and 4) additional SLO and memory experiments in Appendix A.
@@
Figure 5: SLO attainment ($\uparrow$) for end-to-end inference across multiple models and image counts per request. Subfigures (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively. The top and bottom rows show results for 2 and 4 images per request. EPD consistently outperforms all baselines across configurations.
Baselines: We compared the proposed EPD method against two popular baselines: DistServe (Zhong et al., 2024) and vLLM (Kwon et al., 2023). The DistServe baseline implements the prefill-decode (PD) disaggregation approach, in which the prefill and encoding phases are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs. Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying its block manager to accommodate multimodal tokens. The vLLM baseline adopts a monolithic architecture, where all three stages run on the same set of GPUs.
Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 (Yao et al., 2024), InternVL2-8B, and InternVL2-26B (Chen et al., 2024). These LMMs are renowned for their advanced capabilities in processing and understanding multimodal data. Detailed descriptions of the LMMs and their sizes can be found in Appendix E.2.
Datasets: To evaluate performance across diverse scenarios, we use three datasets: synthetic workload, NextQA, and Video-MME. The synthetic workload enables configurable parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings. Unless otherwise noted, the input prompt length is set to 22 tokens. NextQA (Xiao et al., 2021), a benchmark video question-answering dataset, features human-annotated questions and answers, offering a more realistic reflection of real-world video request distributions compared to the synthetic workload. Video-MME (Fu et al., 2024) is a multimodal evaluation dataset designed for assessing large LMMs on video understanding tasks. It contains multiple-choice video question answering items that span diverse video lengths, topics, and reasoning types. In contrast to NextQA, which emphasizes open-ended video-text generation, Video-MME focuses strictly on multiple-choice QA and covers a broader spectrum of video durations and formats.
Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The performance metrics are as follows:
@@
  • TPOT: The average time interval between consecutive output tokens (excluding the first token).
@@
  • Goodput: The highest request rate at which 90% or more SLO attainment is achieved.
For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory in each approach. Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes (see Section 4.3).